OpenStudy (anonymous):

Which of the following points lie in the solution set to the following system of inequalities? y x - 5 y -x - 4 (-5, 2) (5, -2) (-5, -2) (5, 2)

OpenStudy (mertsj):

What does y x - 5 mean?

OpenStudy (anonymous):

sorry lol the first one is the less then or equal too sign and the second one is the more than or equal too sign

OpenStudy (mertsj):

But one has to actually know where these signs are!!

OpenStudy (anonymous):

\[\le then \ge \]

OpenStudy (anonymous):

one minute

OpenStudy (anonymous):

\[y \le x-5 \] \[y \ge -x-4\]

OpenStudy (anonymous):

yea! ^^

OpenStudy (mertsj):

Now plug in the suggested answers and see which makes both sentences true.

OpenStudy (mertsj):

In the ordered pair (-5,2) what is the x value?

OpenStudy (anonymous):

Up above was your same problem and they explained it clearly.

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

No prob. :)